OpenStudy (anonymous):

Three lightships flash simultaneously at 6 a.m. The first flashes every 12 seconds, the second flashes every 30 seconds ant the third flashes every 55 seconds. At what time will the three lightships flash together?

OpenStudy (mathmale):

It seems to me that we should be thinking "smallest common multiple" here. In other words, what is the smallest number divisible by 12, 30 and 55 without remainder(s)? Hint: 12=1*3*2^2, 30=1*2*3*5, and 55=1*5*11.

OpenStudy (anonymous):

you'll need to find the lowest common multiple  therefore 2x2x3x5x11=?